# Beyond Specialization: Assessing the Capabilities of MLLMs in Age and Gender Estimation

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have recently gained immense popularity. Powerful commercial models like ChatGPT-4V and Gemini, as well as open-source ones such as LLaVA, are essentially general-purpose models and are applied to solve a wide variety of tasks, including those in computer vision. These neural networks possess such strong general knowledge and reasoning abilities that they have proven capable of working even on tasks for which they were not specifically trained. We compared the capabilities of the most powerful MLLMs to date: ShareGPT4V, ChatGPT, LLaVA-Next in a specialized task of age and gender estimation with our state-of-the-art specialized model, MiVOLO. We also updated MiVOLO and provide details and new metrics in this article. This comparison has yielded some interesting results and insights about the strengths and weaknesses of the participating models. Furthermore, we attempted various ways to fine-tune the ShareGPT4V model for this specific task, aiming to achieve state-of-the-art results in this particular challenge. Although such a model would not be practical in production, as it is incredibly expensive compared to a specialized model like MiVOLO, it could be very useful in some tasks, like data annotation.
@@ 1 Introduction
Figure 1: An example of evaluated models predictions. The image illustrates output of specialized model (MiVOLO) and different MLLMs. MiVOLO makes predictions based on the face and body crops. Other models make predictions based on prompt and image of body crop.
The rapid development of large multimodal language models (MLLMs or LMMs) has been noteworthy, particularly those integrating language and vision modalities (LVMs). Their advancement is attributed to their high accuracy, generalization capability, reasoning skills, and robust performance with out-of-distribution data. These versatile models excel not only as AI assistants but also in handling unforeseen tasks beyond their initial training scope. The impact of MLLMs is profound, evolving so swiftly that it raises questions about the relevance of specialized models in certain areas. Moreover, there is an increasing interest in using MLLMs for specific computer vision tasks, such as object segmentation, and incorporating them into complex pipelines, such as instruction-based image editing.
In July 2023, we introduced the state-of-the-art model for gender and age determination, MiVOLO [13]. While developing its second version, we explored how competitive MLLMs are in this domain. Initially, ChatGPT4V [28] could not process images of individuals, but this capability was enabled during our project, prompting us to conduct preliminary tests. The results were highly encouraging, leading to a comprehensive evaluation of these neural networks’ potential, including leading open-source solutions like LLaVA [24] [22] and ShareGPT4V [5], which is also based on LLaVA.
@@
For the experiments, we tried to measure all SOTA MLLM models: LLaVA 1.5 and LLaVA-NeXT, ShareGPT4V, and ChatGPT4V. We were unable to measure the newly released Gemini Ultra, as it outright refused to work with images of people.
Since these experiments were conducted as part of improving MiVOLO and comparisons are mostly made with it, we also briefly present the main changes of the new version in this article.
@@
Figure 2: Graphical representation of the study focused on comparison specialized models for age and gender estimation and multimodal models.
@@ 2 Related Works
Neural networks based on transformers with an attention mechanism [41], commonly employed for various natural language processing (NLP) tasks, are also widely utilized for CV tasks. Authors of [42, 11] utilize not only CNNs, but also attention mechanisms, directing the model to focus on features relevant for age estimation. [42] selects the most informative age-specific patches for age estimation. [11] uses transformer to aggregate the sequence of embeddings extracted by CNN and further utilyze the aggregated feature vector for age estimation. MiVOLO [13] employs a transformer to estimate age and gender using face and body crops as input data. In this paper, we introduce MiVOLOv2, which outperforms all specialized models mentioned above.
Model Train Dataset Test Dataset Age MAPE, % $\downarrow$ Age MAE $\downarrow$ Age CS@5, % $\uparrow$ Gender Acc, % $\uparrow$
FP-Age [20] IMDB-clean IMDB-clean - 4.68 63.78 -
MiVOLOv1 LAGENDA IMDB-clean 11.43 [19.25] 4.09 [6.66] 69.72 [47.53] 99.55 [96.74]
224x224 [13] train LAGENDA 13.19 [28.88] 3.99 [7.41] 71.27 [49.64] 97.36 [93.57]
MiVOLOv2 LAGENDA IMDB-clean 11.01 [17.40] 3.97 [6.03] 71.16 [52.11] 99.68 [97.56]
384x384 train LAGENDA 12.06 [24.64] 3.65 [6.16] 74.48 [55.90] 97.99 [95.69]
Table 1: Comparison of MiVOLOv1 and MiVOLOv2 Performance. Both models are evaluated in two modes: $face+body$ and $body\ only$. The results for the $body\ only$ mode are shown in brackets.
@@
3 MiVOLOv2
In this section, we briefly describe the changes in the new version of MiVOLO that we use in subsequent experiments to compare with MLLM models as an anchor specialized model.
@@ 3.1 Evaluation metrics
We utilize the same evaluation metrics as those referenced in previous study [13]:
@@
Additionally, we have slightly modified the Mean Average Percentage Error (MAPE) in this study as follows:
@@
We chose $\epsilon=1$ in the denominator for two main reasons: to prevent division by zero and to mitigate excessively high percentage errors in cases involving infants. For instance, employing $\epsilon=0.083$ (approximately one month) would result in disproportionately large errors for infants, thereby significantly biasing the MAPE.
@@ 3.2 Datasets
We use the same data as before, additionally extending the train dataset by about 40%, so it now contains more than 807,694 samples: 390,730 male and 416,964 female. To do so, we primarily used our production pipelines and some open-source data, like LAION-5B [36]. We selected mostly images where MiVOLOv1 made significant mistakes. We also balanced the right tail of the distribution as much as possible – previously, the training part still suffered from imbalance after about 70 years, despite our efforts to compensate for this.
@@ 3.3 Experiment Details
The training procedure did not change significantly. We conducted many experiments with additional training stage image augmentations but retained only one new augmentation – face blurring, to imitate social network filters or effects from smartphone cameras. In the first stage of training, where we train a single-input model that uses only faces, we apply this blur only slightly, with a 5% random chance. In the second stage, with double-input, we apply blurring with significantly increased parameters, up to almost complete blurring of the face, and with a 70% probability. We also introduce a model trained with new input size 384x384 instead of previously used 224x224. This shows much better results for the in-the-wild domain. For this new model, we decrease dropout and drop-path rates to only 0.1, because the dataset is now large and diverse enough, so less noise caused by dropout works better.
3.4 Results
We briefly present the results in Table 1, which show significant improvements. The results for MiVOLOv2 establish new state-of-the-art results for specialized models. For comparisons with MLLM models, see the following section. Table 2 provides a comparison of results using the Cross-Age Celebrity Dataset (CACD2000) [4] test split.
@@ 3.4 Results
MiVOLOv1 224 [13] 4.09 70.73
MiVOLOv2 384 3.89 73.26
Table 2: Comparison of models using the CACD test split. Both versions of MiVOLO are evaluated in face + body mode.
@@ 4.1 Benchmarks
  • IMDB-clean dataset [21] used in the MiVOLO article [13]. The dataset was enhanced with body bounding boxes associated with facial pairs. It comprises 183,886 training images, 45,971 validation images, and 56,086 test images. We use only the test images for our evaluation.
  • LAGENDA [13], a dataset well-balanced in terms of age and gender attributes, features face-body pairs with ground truth obtained through human annotations via weighted voting. It includes proprietary training and validation parts, and an open-source test part, containing 67,159 images from the Open Images Dataset [14] featuring 84,192 individuals aged from 0 to 95.
@@
Initially, the OpenAI API imposed a limit of 100 requests per day for the gpt-4-vision-preview models, which has since been increased to 1,500 requests per day. Due to this limitation, we randomly selected a small subset from LAGENDA to run with ChatGPT4V and included it in our comparison. We selected 200 random samples for each age group at intervals of 5 years (e.g., 0-5, 5-10, etc.) for ages $[0;90]$. However, $>21\%$ of these images had to be removed because ChatGPT refused to provide answers. As a result, this dataset contains 3,062 samples. We refer to this dataset as NanoLAGENDA.
We opted for LAGENDA over IMDB to minimize the risk that MLLMs would provide correct answers not through age and gender estimation but because of its familiarity with famous individuals, well-known movies, etc. On the other hand, LAGENDA, annotated by human annotators, does not have actual ground truths for labels. Nevertheless, we chose it because the risk of this drawback is lower. We will demonstrate later in this article that our assumption of LAGENDA’s labels being very close to actual ground truths on average, based on measurements from [13], is valid.
@@ 4.2 Method
Base Model Body or Face Crop Age MAE $\downarrow$ Age CS@5, % $\uparrow$ Gender Acc, % $\uparrow$
@@
Prompt for MLLMs
@@
We set the temperature for all models to 0.0. For ChatGPT, we additionally set the parameter ”seed” to 1234 and ”n” to 1. The latter is necessary due to reports that just the seed and temperature are not sufficient to ensure deterministic and reproducible results for vision model via API.
@@
Model Input Age Acc, % $\uparrow$ MAE, % $\downarrow$ Gender Acc, % $\uparrow$
OridinalCLIP [19] Face 61.2 ± 4.2 0.47 ± 0.6 -
L2RCLIP [43] Face 66.2 ± 4.4 0.36 ± 0.5 -
MiVOLOv1 224x224 [13] Body & face 68.68 0.345 96.5
MiVOLOv2 384x384 Body & face 69.43 0.333 97.39
ShareGPT4V 7B fine-tuned Body 66.7 0.349 95.65
ShareGPT4V 7B fine-tuned with faces Face 67.95 0.338 96.63
Table 4: Comparison of models using Adience benchmark. MAE here is calculated on classification labels.
Model Input Age MAPE, % $\downarrow$ Age MAE $\downarrow$ Age CS@5, % $\uparrow$ Gender Acc, % $\uparrow$
LLaVA 1.5 7B [24] Entire body 16.86 7.59 $\pm$ 7.22 48.79 99.38
Body w/o face 43.18 18.44 $\pm$ 17.38 25.71 94.85
LLaVA-NeXT 34B[23] Entire body 13.73 6.20 $\pm$ 5.78 55.40 99.51
Body w/o face 23.03 9.58 $\pm$ 8.51 39.67 97.82
ShareGPT4V 7B [5] Entire body 16.71 7.16 $\pm$ 6.48 53.24 99.44
Body w/o face 25.80 11.16 $\pm$ 10.45 39.07 97.24
ChatGPT4V [29] Entire body 12.12 4.66 $\pm$ 4.51 68.10 98.43
Body w/o face 22.56 7.82 $\pm$ 7.52 49.15 93.02
MiVOLOv2 384x384 Body & face 11.61 4.33 $\pm$ 4.47 69.90 97.71
Body w/o face 21.82 7.19 $\pm$ 7.67 49.14 95.61
ShareGPT4V 7B fine-tuned Entire body 10.95 4.22 $\pm$ 3.99 72.09 99.51
Body w/o face 20.53 7.64 $\pm$ 7.35 49.95 97.48
ShareGPT4V 13B fine-tuned Entire body 11.30 4.26 $\pm$ 4.18 73.34 99.44
Body w/o face 19.90 7.40 $\pm$ 7.72 54.01 98.10
@@ 4.3 LLaVA finetune
Building on insights from the MiVOLO paper, simultaneous training for both tasks has shown to be advantageous. Consequently, our training approach mirrors the evaluation methodology described in the preceding section.
@@
The optimal hyperparameters identified through our experiments are as follows: learning rate = $2e^{-6}$, LoRA disabled (full fine-tuning), per_device_train_batch_size = 32, num_train_epochs = 1, checkpoint every 450 iterations, warmup_ratio = 0.03, lr_scheduler_type ”cosine”
@@
Future developments could explore dual-crop inputs (body + face as separate images), as seen in the original MiVOLO model. However, the feasibility of training existing models with multiple images per input remains an open question.

### v4

@@ Abstract
Multimodal Large Language Models (MLLMs) have recently gained immense popularity. Powerful commercial models like ChatGPT and Gemini, as well as open-source ones such as LLaVA, are essentially general-purpose models and are applied to solve a wide variety of tasks, including those in computer vision. These neural networks possess such strong general knowledge and reasoning abilities that they have proven capable of working even on tasks for which they were not specifically trained. We compared the capabilities of the most powerful MLLMs to date including ShareGPT4V, ChatGPT 4V/4O, and LLaVA Next in the specialized task of age and gender estimation, with the state-of-the-art specialized model MiVOLO384. In our study, we discovered that the fine-tuned open-source ShareGPT4V model is capable of outperforming the specialized model in age and gender estimation tasks. At the same time, the proprietary ChatGPT-4O beats both in the age estimation task but does not perform as confidently in gender recognition. This gives interesting insights about the strengths and weaknesses of the participating models and suggests that with a few tweaks, general-purpose MLLM models can match or even surpass specialized ones in certain fields. Even though these fine-tuned models might require more computing power, they offer big benefits for tasks where computing power is not a limiting factor and where the best accuracy is key, such as data annotation.
Keywords:
MLLM, VLM, Human Attribute Recognition, Age estimation, Gender estimation, Large Models Generalization
@@ 1 Introduction
Figure 1: An example of evaluated models predictions. The image illustrates output of specialized $MiVOLO_{384}$ model and different MLLMs. $MiVOLO_{384}$ makes predictions based on the face and body crops. Other models make predictions based on prompt and image of body crop.
The rapid development of multimodal large language models (MLLMs or LMMs) has been noteworthy, particularly those integrating language and vision modalities (LVMs). Their advancement is attributed to their high accuracy, generalization capability, reasoning skills, and robust performance with out-of-distribution data. These versatile models excel not only as AI assistants but also in handling unforeseen tasks beyond their initial training scope. The impact of MLLMs is profound, evolving so swiftly that it raises questions about the relevance of specialized models in certain areas. Moreover, there is an increasing interest in using MLLMs for specific computer vision tasks, such as object segmentation, and incorporating them into complex pipelines, such as instruction-based image editing.
We explored the competitiveness of MLLMs in the specific domain of age and gender estimation. Initially, we conducted preliminary tests with ChatGPT-4V [29]. The results were highly encouraging, prompting a comprehensive evaluation of these neural networks’ potential, including leading open-source solutions such as LLaVA [25, 23] and ShareGPT-4V [5], which is also based on LLaVA. Later, we updated this work with results for ChatGPT-4O, the newest and most powerful OpenAI model available at the time.
@@
For the experiments, we tried to measure SOTA MLLM models: LLaVA 1.5 and LLaVA-NeXT, ShareGPT4V, and ChatGPT4. We were unable to measure the newly released Gemini Ultra, as it outright refused to work with images of people.
We’ve also made improvements to the state-of-the-art specialized model MiVOLO [14] to ensure fair competition among cutting-edge models.
@@
Figure 2: Graphical representation of the study focused on comparison specialized model for age and gender estimation and multimodal models.
@@ 2 Related Works
Neural networks based on transformers with an attention mechanism [42], commonly employed for various natural language processing (NLP) tasks, are also widely utilized for CV tasks. Authors of [43, 11] utilize not only CNNs, but also attention mechanisms, directing the model to focus on features relevant for age estimation. [43] selects the most informative age-specific patches for age estimation. [11] uses transformer to aggregate the sequence of embeddings extracted by CNN and further utilyze the aggregated feature vector for age estimation. MiVOLO [14] employs a transformer to estimate age and gender using face and body crops as input data. In this paper, we enhance MiVOLO, resulting in a model that outperforms all the specialized models mentioned above.
@@
3 Enhanced MiVOLO Model
In this section, the improvements made to the MiVOLO [14] model are briefly described to achieve state-of-the-art performance. The enhanced model is used in subsequent experiments as a benchmark to compare with MLLMs, serving as an anchor specialized model. The original model from [14] is referred to as $MiVOLO_{224}$, as it was trained with 224x224 input size. We train the MiVOLO model with 384x384 input size on the extended train dataset. The enhanced model is referred to as $MiVOLO_{384}$ accordingly.
@@ 3.1 Evaluation Metrics
The same evaluation metrics as in the original study [14] were utilized:
@@
Additionally, the Mean Average Percentage Error (MAPE) was slightly modified in this study as follows:
@@
An $\epsilon=1$ was chosen in the denominator for two main reasons: to prevent division by zero and to mitigate excessively high percentage errors in cases involving infants. For instance, employing $\epsilon=0.083$ (approximately one month) would result in disproportionately large errors for infants, thereby significantly biasing the MAPE.
@@ 3.2 Datasets
The same data as in the MiVOLO paper [14] was used, with an extension of the training dataset by approximately 40%, resulting in over 807,694 samples: 390,730 images of males and 416,964 images of females. The extended version of LAGENDA[14] train dataset is referred to as $LAGENDA_{ext}$. The extension was achieved primarily through production pipelines and supplemented with open-source data, such as LAION-5B [37]. Focus was given to selecting images where the original $MiVOLO_{224}$ model made significant errors. Additionally, efforts were made to balance the distribution’s right tail, as the original training dataset was imbalanced for ages above 70 years. The test LAGENDA dataset was taken unchanged.
@@ 3.3 Experiment Details
Many experiments were conducted with additional training stage image augmentations, but only one new augmentation face blurring was retained, to imitate social network filters or effects from smartphone cameras. In the first stage of training, where a single-input model that uses only faces was trained, this blur was applied slightly, with a 5% random chance. In the second stage, with double-input, blurring parameters were significantly increased, with up to a 70% probability. The model was also trained with a 384x384 input size instead of the originally used 224x224, showing much better results for the in-the-wild domain. Dropout and drop-path rates were decreased to 0.1, due to the large and diverse dataset.
Model Train Dataset Test Dataset MAPE, % $\downarrow$ MAE $\downarrow$ CS@5, % $\uparrow$
FP-Age [21] IMDB-clean IMDB-clean - 4.68 63.78
$MiVOLO_{224}$[14] LAGENDA IMDB-clean 11.43 4.09 69.72
LAGENDA 13.19 3.99 71.27
$MiVOLO_{384}$ $LAGENDA_{ext}$ IMDB-clean 11.01 3.97 71.16
LAGENDA 12.06 3.65 74.48
Table 1: Comparison of $MiVOLO_{224}$ and $MiVOLO_{384}$. Age performance in face + body mode.
Model Train Dataset Test Dataset MAPE, % $\downarrow$ MAE $\downarrow$ CS@5, % $\uparrow$
$MiVOLO_{224}$[14] LAGENDA IMDB-clean 19.25 6.66 47.53
LAGENDA 28.88 7.41 49.64
$MiVOLO_{384}$ $LAGENDA_{ext}$ IMDB-clean 17.40 6.03 52.11
LAGENDA 24.64 6.16 55.90
Table 2: Comparison of $MiVOLO_{224}$ and $MiVOLO_{384}$. Age performance in body only mode. LAGENDA in the train column refers to a part of the dataset used for training, as described in [14].
Model Train Dataset Test Dataset Gender Acc, % $\uparrow$
FP-Age [21] IMDB-clean IMDB-clean -
$MiVOLO_{224}$[14] LAGENDA IMDB-clean 99.55
LAGENDA 97.36
$MiVOLO_{384}$ $LAGENDA_{ext}$ IMDB-clean 99.68
LAGENDA 97.99
Table 3: Comparison of $MiVOLO_{224}$ and $MiVOLO_{384}$. Gender Accuracy.
@@
$MiVOLO_{224}$[14] [14] 4.09 70.73
$MiVOLO_{384}$ 3.89 73.26
Table 4: Comparison of models using the CACD test split. MiVOLO models are evaluated in face + body mode.
Model Age Acc, % $\uparrow$ Gender Acc, % $\uparrow$
$MiVOLO_{224}$ [14] 61.07 95.73
$MiVOLO_{384}$ 62.28 97.5
Table 5: Comparison of models using the FairFace validation margin125 split. MiVOLO models are evaluated in face + body mode.
3.4 Results
Tables 1, 2 show comparison of original $MiVOLO_{224}$ and $MiVOLO_{384}$ models. The results for $MiVOLO_{384}$ establish new state-of-the-art results for specialized models. For comparisons with MLLM models, see the following section. Tables 4, 5 provide a comparison of results using the Cross-Age Celebrity Dataset (CACD2000) [4] test split and FairFace [12] validation ‘margin125‘ split.
@@ 3.4 Results
In this section, we compare the capabilities of MLLMs with MiVOLO in age and gender estimation tasks on various benchmarks. Additionally, we investigate the effects of fine-tuning MLLMs on large target datasets to enhance their accuracy in these specific tasks.
@@ 4.1 Benchmarks
  • IMDB-clean dataset [22], which was used in the MiVOLO article [14]. The original IMDB-clean dataset was enhanced with body bounding boxes associated with facial pairs. It comprises 183,886 training images, 45,971 validation images, and 56,086 test images. We use only the test images for our evaluation.
  • LAGENDA benchmark [14], a dataset well-balanced in terms of age and gender attributes, features face-body pairs with ground truth obtained through human annotations via weighted voting. It includes proprietary training and validation parts, and an open-source test part, containing 67,159 images from the Open Images Dataset [15] featuring 84,192 individuals aged from 0 to 95.
@@
Initially, the OpenAI API imposed a limit of 100 requests per day for the gpt-4-vision-preview models, which has since been increased to 1,500 requests per day. Due to this limitation, we randomly selected a small subset from LAGENDA [14] to run with ChatGPT4V and included it in our comparison. We selected 200 random samples for each age group at intervals of 5 years (e.g., 0-5, 5-10, etc.) for ages $[0;90]$. However, $>21\%$ of these images had to be removed because ChatGPT refused to provide answers. As a result, this dataset contains 3,062 samples. We refer to this dataset as NanoLAGENDA.
We opted for LAGENDA over IMDB to minimize the risk that MLLMs would provide correct answers not through age and gender estimation but because of its familiarity with famous individuals, well-known movies, etc. On the other hand, LAGENDA, annotated by human annotators, does not have actual ground truths for labels. Nevertheless, we chose it because the risk of this drawback is lower.
@@ 4.2 Method
Base Model Body or Face Crop Age Gender
MAE $\downarrow$ CS@5, % $\uparrow$ Acc, % $\uparrow$
@@
Prompt for MLLMs Please provide the perceived gender and estimated age of the individual in this photo in the format ’[gender; age]’, where the age must be a single integer between 0 and 130. It should not be a range. The gender must be one of two options: ’male’ or ’female’. I understand that your response will be based on visual analysis and may not accurately reflect the actual age or gender identity of the individual.
@@
We set the temperature for all models to 0.0. For ChatGPT, we additionally set the parameter seed to 1234 and n to 1. The latter is necessary due to reports that just the seed and temperature are not sufficient to ensure deterministic and reproducible results for vision model via API.
@@
Model Input Age Acc, % $\uparrow$ Age MAE, % $\downarrow$ Gender Acc, % $\uparrow$
OridinalCLIP [20] Face 61.2 0.47 -
L2RCLIP [44] Face 66.2 0.36 -
$MiVOLO_{224}$ [14] Body & face 68.68 0.345 96.5
$MiVOLO_{384}$ Body & face 69.43 0.333 97.39
ShareGPT4V 7B FT Body 66.7 0.349 95.65
ShareGPT4V 7B FT Face 67.95 0.338 96.63
Table 7: Comparison of models using Adience benchmark. MAE here is calculated on classification labels. FT denotes models that are fine-tuned version on the corresponding input.
Model Input Age Gender
MAPE, % $\downarrow$ MAE $\downarrow$ CS@5, % $\uparrow$ Acc, % $\uparrow$
LLaVA 1.5 7B [25] Entire body 16.86 7.59 48.79 99.38
W/o face 43.18 18.44 25.71 94.85
LLaVA-NeXT 34B[24] Entire body 13.73 6.20 55.40 99.51
W/o face 23.03 9.58 39.67 97.82
ShareGPT4V 7B [5] Entire body 16.71 7.16 53.24 99.44
W/o face 25.80 11.16 39.07 97.24
ChatGPT4V [30] Entire body 12.12 4.66 68.10 98.43
W/o face 22.56 7.82 49.15 93.02
ChatGPT4O [30] Entire body 10.42 4.07 73.91 98.66
W/o face 15.39 5.73 60.17 96.92
$MiVOLO_{384}$ Body & face 11.61 4.33 69.90 97.71
W/o face 21.82 7.19 49.14 95.61
ShareGPT4V 7B FT Entire body 10.95 4.22 72.09 99.51
W/o face 20.53 7.64 49.95 97.48
ShareGPT4V 13B FT Entire body 11.30 4.26 73.34 99.44
W/o face 19.90 7.40 54.01 98.10
@@ 4.3 LLaVA Finetune
Building on insights from the MiVOLO [14], simultaneous training for both tasks has shown to be advantageous. Consequently, our training approach mirrors the evaluation methodology described in the preceding section.
@@
The optimal hyperparameters identified through our experiments are as follows:
Hyperparameter Value
Learning rate $2\times 10^{-6}$
LoRA disabled (full fine-tuning)
Per-device train batch size 32
Number of train epochs 1
Checkpoint every 450 iterations
Warmup ratio 0.03
LR scheduler type cosine
@@
Future developments could explore dual-crop inputs (body + face as separate images), as seen in the original $MiVOLO_{224}$ model. However, the feasibility of training existing models with multiple images per input remains an open question.